# The emergent dynamics of double-folded randomly branching ring polymers

### v1

@@ IV.3 Scaling analysis of the Emergent Dynamics
Before analyzing in detail the speed-up associated with the combined dynamics, we first present a detailed scaling analysis of the dynamics reported in Fig. 4. This analysis begins with the three limiting cases discussed in the Theoretical Background (Section II) and concludes with the mixed dynamics.
@@
Specifically, in Fig. 5 we report scaling results for $g_{1}$ and $g_{3}$ in the context of the exclusive application of repton moves (slithering dynamics), for ideal trees. In Fig. 6, we report corresponding scaling results in the context of three other dynamics – the RiAO dyanmics (left column), the Brownian motion of trees (central column) and the combined dynamics (right column) – for ideal rings (top row), self-avoiding rings (middle row), and rings in melts (bottom row). Additional details and figures can be found in Section II of the Supplementary Material.
IV.3.1 Slithering dynamics around fixed ideal trees
@@
Let us first consider the slithering dynamics around a fixed (quenched) tree emerging from the exclusive application of repton moves. In Fig. 5 in the Supplementary Material, we show the corresponding monomer and CM MSD for ideal randomly branching double-folded rings composed of $N_{m}=64,216,512,1000$ monomers. An obvious feature is the nearly absent CM motion of the rings, $g_{3}(t)<\langle R_{g}^{2}(N_{m})\rangle/N_{m}$, such that $g_{1}(t)\approx g_{2}(t)$. We then use various rescaling of the same data in the different panels of the figure to reveal different time and length scales relevant to the monomer and the CM motion. Namely, the ring monomers are fully mobile while remaining bound to the quenched tree conformations that the rings are wrapping. As a consequence, $\lim_{t\rightarrow\infty}g_{2}(t)=2\langle R_{g}^{2}(N_{m})\rangle$. The asymptotic plateau is reached on the reptation time scale, $\tau_{rep}\sim N_{m}^{3}$, of the rings, while the crossover from the $t^{\nu/2}$ to the $t^{\nu}$ regime in monomer diffusion occurs at their Rouse time, $\tau_{R}\sim N_{m}^{2}$.
@@
Figure 5: Scaling analysis of the slithering dynamics of ideal randomly branching double-folded rings emerging from the exclusive application of repton moves in our model.
@@
Figure 6: Scaling behavior for $g_{1}(t)$ (diamonds) and $g_{3}(t)$ (circles) for three systems: ideal (top row), isolated self-avoiding (middle row), and double-folded ring polymers in melts (bottom row). Time is rescaled by the relaxation time, defined as $\tau_{\text{max}}=\frac{\langle R_{g}^{2}\rangle}{D}$. Left column: Results for RiAO-like dynamics. Middle column: Results for Brownian dynamics of randomly quenched trees. Right column: Results for a mixed of the two dynamics, with $p_{m}=0.5$ in Eq. 10.
IV.3.2 RiAO-like dynamics
@@
As already discussed in detail in Ref. [24], the emerging dynamics resulting from a combination of repton and hairpin moves agrees with the theoretical predictions for the RiAO model (Refs. [2, 6] and Sec. II). As shown in the left column of Fig. 6, an appropriate rescaling of our present data leads again to good agreement with the predicted $g_{1}(t)\sim t^{\frac{2\nu}{2+\rho}}t$ and $g_{3}(t)\sim t^{\frac{2\nu+1}{2+\rho}}$ for the monomer and tree CM dynamics. The corresponding effective dynamic exponents are shown in the insets of the panels. At least qualitatively, their slow decay in the tree scaling regime agrees with what one would expect from estimating the dynamic exponents from the effective static exponents, $\rho_{eff}$ and $\nu_{eff}$, shown in Fig. 3. As shown in the middle column of Fig. 4 (green crosses), there is also good agreement between the predicted and the observed ring weight dependence of the asymptotic CM diffusion constants.
IV.3.3 Brownian dynamics of randomly quenched trees
@@
The emerging dynamics resulting from the exclusive application of tree node moves to our double-folded rings is presented in the central column of Fig. 6.
@@
For ideal and self-avoiding trees, our data are in good agreement with the predictions of the Rouse model (Tabel 1). The ring CM essentially exhibits simple diffusion, $g_{3}(t)\sim t/N_{m}$, with a diffusion constant inversely proportional to the number of ring monomers. A slowdown occurs independently of the ring size only over the first $50$ MC sweeps in the ideal trees, while in the self-avoiding case, the slowdown persists for a longer duration, up to $1000$ sweeps.
The monomer dynamics, $g_{1}(t)$, in the intermediate intra-chain scaling regime also approaches the prediction $g_{1}(t)\sim t^{\frac{2\nu}{2\nu+1}}$ of the Rouse model (Table 1). Again, the slow decay of the effective exponents reported in the detailed analysis in Fig. 6 agrees, at least qualitatively, with the observed crossover of the static exponents in Fig. 3.
@@
Interestingly, the observed monomer MSDs in the melt are also in reasonable accord with the prediction $g_{1}(t)\sim t^{0.4}$ of the Rouse model. However, the ring CM clearly exhibits anomalous diffusion in the intermediate regime. Experimental studies on ring polymers in the melt state have reported a substantial slowdown in COM dynamics  [46, 47, 48], which occurs before the system reaches the free diffusion regime.
IV.3.4 Combining monomer and node moves
@@
Given that the node dynamics is significantly faster than the monomer dynamics in ideal and isolated self-avoiding systems (see Fig. 4), it is unsurprising that the former dominates the combined dynamics entirely, causing the resulting scaling plots to essentially collapse onto those in the central column.
It is tempting to draw similar conclusions in the melt case, but with respect to the left column of Fig. 6 and a faster RiAO-like dynamics (bottom row in Fig. 4). Yet, the situation is more subtle since the overal dynamics is accelerated with respect to the pure dynamics of either monomer moves or tree node moves. The excellent scaling of the data reported in the bottom right panel of Fig. 6 is thus in itself an interesting observation. We have tentatively included slopes predicted by the FLG model, but it is difficult to draw any quantiative conclusions at this point.
V Discussion
@@
The remaining challenge is to develop a protocol for the analysis of data obtained in mixed dynamical schemes. In particular, this protocol should allows us to identify and distinguish situations where the contributions of different local modes to the emergent dynamics are independent from those where the emergent dynamics arise from coupling effects.
Figure 7: Left column: The asymptotic center-of-mass (CM) diffusion constants for different systems, obtained using a combination of monomer and node moves in varying proportions ($0\leq p_{m}\leq 1$), as described in Eq. 10. The dashed lines represent the predictions from the null model (Eq. 11). Right column: The difference between the observed asymptotic CM diffusion constants and the expected values from the null model, as $\Delta D(p_{m})=D(p_{m})-\left(p_{m}D_{\mathrm{RiAO}}+(1-p_{m})D_{\mathrm{QTree}}\right)$. For better visualization across different system sizes, all data are normalized by $D_{\text{QTree}}$.
@@
In Appendix A we propose, with Eq. (20), an approximate plausible generalization of Eq. (11) for the full time dependence of the null model, for the monomer and CM MSDs in a mixed scheme. In Fig. 9 in Appendix B, we show that Eq. (20) works well in a case where we expect the null model to apply: the dynamics emerging from the combination of repton and tree node moves, which are indeed completely decoupled within our model.
@@
While the ergodic RiAO dynamics provides an interesting example of the emergence of a qualitatively different dynamics from the combination of local moves (see Appendix C), in the remainder of the (main) text, we focus on the emerging dynamics resulting from the combination of monomer and node moves in different proportions, $0\leq p_{m}\leq 1$, see Eq. 10. As evident from Figs. 4 and 10, in interacting systems, this combination can lead to an accelerated dynamics compared to the two limiting cases: the RiAO dynamics ($p_{m}=1$) and the dynamics of quenched trees ($p_{m}=0$). This raises the question of how this acceleration depends on the ring weight, $N_{m}$, and the mixing parameter, $p_{m}$.
@@
For our mixing rule, Eq. (10), of the monomer Supplemantaryand node moves, the prediction of the null model, Eq. (11), reads $D(p_{m})=p_{m}D_{RiAO}+(1-p_{m})D_{\mathrm{QTree}}$ or, equivalently, $D(p_{m})/D_{\mathrm{QTree}}=1+(D_{RiAO}/D_{\mathrm{QTree}}-1)p_{m}$. The panels in the left column of Fig. 7 show the corresponding results for the three types of investigated systems (ideal rings, self-avoided rings and rinds in melts) – note that the normalization to $D_{\mathrm{QTree}}$ allows us to show results for rings of different sizes on a common linear scale (Raw results of $D$ are presented in Fig. LABEL:fig_SM:Diffusion_tau_raw in the Supplemantary Material.).
@@
Results reveal the importance and character of deviations of the emergent dynamics from the null model. Namely, while the emergent dynamics in the melt is significantly accelerated and dominated by the deviations from the null model, the corrections appear to be more of a qualitative nature for self-avoiding rings and statistically insignificant for ideal rings. This said, our decision to normalize the diffusion constants in all three cases by those of the asymptotically faster Brownian Tree Dynamics (right column of Fig. 4) can lead to the wrong impression that $D(p_{m})$ extrapolates to zero for finite ring sizes (as it does indeed in the case of the slithering dynamics in Fig. 9(d)). However, in the present case $D(p_{m})/D_{\mathrm{QTree}}$ remains finite for all $N_{m}$, albeit much smaller than unity for ideal and self-avoiding double-folded rings and equal to zero only in the limit of infinite chain length.
@@
In the right column of Fig. 7, we show the difference $\Delta D(p_{m})=D(p_{m})-\left(p_{m}D_{\mathrm{RiAO}}+(1-p_{m})D_{\mathrm{QTree}}\right)$ between the observed asymptotic CM diffusion constant and the expected value from the null model. To simplify a comparison with the data shown in the panels in the left column, we do not show absolute values for $\Delta D(p_{m})$ but $\Delta D(p_{m})/D_{\mathrm{QTree}}$. If $\Delta D(p_{m})$ vanishes by construction in the two limits $p_{m}=0$ and $p_{m}=1$ (i.e., a value of zero is in this case a “true” and not just an apparent zero), our data suggest an intriguing symmetry in how the node and the monomer dynamics combine.
The most general admissible expression for contributions $\Delta D_{AB}$ to the ring CM diffusion constant, $D$, arising from a coupling of two distinct $A$- and $B$-types of dynamics would seem to be a sum of the form
$\Delta D_{AB}(p_{A},p_{B})$
$=$ $\sum_{i}a_{i}\left(\left(p_{A}D_{A}\right)^{\alpha_{i}}\left(p_{B}D_{B}\right)^{1-\alpha_{i}}\right)^{\beta_{i}}\left(p_{A}D_{A}+p_{B}D_{B}\right)^{1-\beta_{i}}$
with exponents $0\leq\alpha_{i}$, assuring that (i) $\Delta D_{AB}\neq 0$ if and only if both types of the elementary dynamics are present, (ii) $\Delta D_{AB}$ has the same units as $D_{A}$ and $D_{B}$, and (iii) the coupled dynamics is slowed down by the same factor as the $A$- and the $B$-dynamics, when both types of moves are rejected with equal probabilities, $p_{A}=p_{B}=p$:
$\Delta D_{AB}(p_{A}=p,p_{B}=p)$
$=$ $\sum_{i}a_{i}\left(\left(pD_{A}\right)^{\alpha_{i}}\left(pD_{B}\right)^{1-\alpha_{i}}\right)^{\beta_{i}}\left(pD_{A}+pD_{B}\right)^{1-\beta_{i}}$
$=$ $p\sum_{i}a_{i}\left(D_{A}^{\alpha_{i}}D_{B}^{1-\alpha_{i}}\right)^{\beta_{i}}\left(D_{A}+D_{B}\right)^{1-\beta_{i}}$
$=$ $p\,\Delta D_{AB}(p_{A}=1,p_{B}=1)$ (14)
@@
Interestingly, our results for $\Delta D_{AB}$ reported in the right column of Fig. 7 suggest that the coupling contribution adopts the simplest possible form of
@@
Is this (in hindsight) obvious and inevitable or, alternatively, does our observation constrain the choice of admissible microscopic coupling mechanisms? Since the CM diffusion of our rings is diffusive in the long time limit, we consider as the simplest model the coupling of two diffusion steps, $\vec{\delta}_{A}$ and $\vec{\delta}_{B}$, for a particle moving under the influence of two random processes $A$ and $B$. To define the additional displacement, $\vec{\delta}_{AB}=\delta_{AB}\hat{\delta}_{AB}$, due to the coupling of the two steps, we need to specify its magnitude $\delta_{AB}\equiv\lVert\vec{\delta}_{AB}\rVert$ and direction, $\hat{\delta}_{AB}\equiv\vec{\delta}_{AB}/\lVert\vec{\delta}_{AB}\rVert$. For the magnitude, admissible combinations of $\delta_{A}$ and $\delta_{B}$ have to obey the same dimensionality constraint, that we have discussed above for the diffusion constant. With simple vector operations like $\lVert\vec{\delta}_{A}\times\vec{\delta}_{B}\rVert$ and $\vec{\delta}_{A}\cdot\vec{\delta}_{B}$ having the unit of the square of a length, one arrives rather naturally at a scaling, $\delta_{AB}\propto\sqrt{\delta_{A}\delta_{B}}$, of the length of the coupling steps with the geometric mean of the elementary step lengths, that may be modulated by geometric factors. On first sight, this seems to directly lead to the observed scaling $\Delta D_{AB}=\langle\delta_{AB}^{2}\rangle\propto\sqrt{\langle\delta_{A}^{2}\rangle\,\langle\delta_{B}^{2}\rangle}\propto\sqrt{D_{A}D_{B}}$ of the coupling contribution to the diffusion constant. However, the relation $\Delta D_{AB}\equiv D_{AB}-D_{A}-D_{B}$ only holds, if additionally $\hat{\delta}_{AB}=\left(\vec{\delta}_{A}\times\vec{\delta}_{B}\right)/\lVert\vec{\delta}_{A}\times\vec{\delta}_{B}\rVert$ or if the direction of the coupling steps, $\hat{\delta}_{AB}$, is not correlated with $\vec{\delta}_{A}$ and $\vec{\delta}_{B}$. In contrast, if $\hat{\delta}_{AB}$ lies in the plane spanned by the elementary steps, $\vec{\delta}_{A}$ and $\vec{\delta}_{B}$, correlations lead to more complicated relations of the general type Eq. (V).
As our final point, Fig. 8 addresses the ring size dependence of the magnitude of $\Delta D_{AB}$ arising from the coupling of the Brownian tree and the ring-in-an-array-of-obstacles-like dynamics. In spite of the vast difference in the amplitude of the Brownian tree dynamics for isolated self-avoiding rings and for rings in a melt (Fig. 4), the absolute magnitude of the CM motion due to the coupling with the RiAO-like dynamics is quite similar for the two systems. Our data for isolated self-avoiding rings are in good agreement with a Rouse-like scaling, $\Delta D_{AB}\sim N^{-1}$. Our data for melts are less clear, but like the diffusion constants measured for pure Brownian tree dynamics, they are compatible with a slow crossover to a Rouse-like scaling. Notably the total CM diffusion constant
@@
in the melt case seems to be dominated asymptotically by the mobility arising from the coupling of the Brownian tree and the RiAO-like dynamics.
VI Summary and Conclusion
Ring polymers pose a number of challenging fundamental problems in polymer physics. In this article, we have studied the possible dynamics of a (lattice) model of tightly double-folded chains [1, 2, 3, 6, 24]. If the long-time goal is to apply the model to specific examples from a wide range of ring polymer systems (see e.g. [8, 12, 13, 14, 15, 16, 17, 18]), our objective here was to study generic aspects of the coupling of different local modes of dynamics. In particular, we have exploited the opportunity inherent in a mesoscale description to freely vary the relative kinetic rates instead of being limited to ratios emerging in particular experimentally or numerically accessible target or model systems.

### v2

@@ 4.3 Scaling analysis of the Emergent Dynamics
Before analyzing in detail the speed-up associated with the combined dynamics, we first present a detailed scaling analysis of the dynamics reported in Fig. 4. This analysis begins with the three limiting cases discussed in the Theoretical Background (Section 2) and concludes with the mixed dynamics.
@@
Specifically, in Fig. 5 we report scaling results for $g_{1}$ and $g_{3}$ in the context of the exclusive application of repton moves (slithering dynamics), for ideal trees. In Fig. 6, we report corresponding scaling results in the context of three other dynamics: the RiAO dyanmics (left column), the Brownian motion of trees (central column) and the combined dynamics (right column), for ideal rings (top row), self-avoiding rings (middle row), and rings in melts (bottom row). Additional details and figures are provided in Figs. S9–S15 of Sections VIII–IX in the Supplementary Material.
4.3.1 Slithering dynamics around fixed ideal trees
@@
Let us first consider the slithering dynamics around a fixed (quenched) tree emerging from the exclusive application of repton moves. In Fig. 5 (see also Fig. S15 in the Supplementary Material for more detailed results), we show the corresponding monomer and CM MSD for ideal randomly branching double-folded rings composed of $N_{m}=64,216,512,1000$ monomers. An obvious feature is the nearly absent CM motion of the rings, $g_{3}(t)<\langle R_{g}^{2}(N_{m})\rangle/N_{m}$, such that $g_{1}(t)\approx g_{2}(t)$. We then use various rescaling of the same data in the different panels of the figure to reveal different time and length scales relevant to the monomer and the CM motion. The results are shown in Fig. 5. Namely, the ring monomers are fully mobile while remaining bound to the quenched tree conformations that the rings are wrapping. As a consequence, $\lim_{t\rightarrow\infty}g_{2}(t)=2\langle R_{g}^{2}(N_{m})\rangle$. The asymptotic plateau is reached on the reptation time scale, $\tau_{rep}\sim N_{m}^{3}$, of the rings, while the crossover from the $t^{\nu/2}$ to the $t^{\nu}$ regime in monomer diffusion occurs at their Rouse time, $\tau_{R}\sim N_{m}^{2}$.
@@
Figure 5: Scaling analysis of the slithering dynamics of ideal, randomly branching, double-folded rings emerging from the exclusive application of repton moves in our model. The inset graphs in each panel share the same x-tick positions as the main panel they accompany.
@@
Figure 6: Scaling behavior for $g_{1}(t)$ (crosses) $g_{2}(t)$ (diamonds) and $g_{3}(t)$ (circles) for three systems: ideal (top row), isolated self-avoiding (middle row), and double-folded ring polymers in melts (bottom row). Time is rescaled by the relaxation time, defined as $\tau_{\text{max}}=\frac{\langle R_{g}^{2}\rangle}{D}$. Left column: Results for RiAO-like dynamics. Middle column: Results for Brownian dynamics of randomly quenched trees. Right column: Results for a mixed of the two dynamics, with $p_{m}=0.5$ in Eq. 10.
4.3.2 RiAO-like dynamics
@@
As already discussed in detail in Ref. 38, the emerging dynamics resulting from a combination of repton and hairpin moves agrees with the theoretical predictions for the RiAO model (Refs. 2, 6 and Sec. 2). As shown in the left column of Fig. 6, an appropriate rescaling of our present data leads again to good agreement with the predicted $g_{1}(t)\sim t^{\frac{2\nu}{2+\rho}}$ and $g_{3}(t)\sim t^{\frac{2\nu+1}{2+\rho}}$ for the monomer and tree CM dynamics (Detailed analysis is provided in Figs. S12–S14 in Sec. IX of the Supplementary Material). The corresponding effective dynamic exponents are shown in the insets of the panels. At least qualitatively, their slow decay in the tree scaling regime agrees with what one would expect from estimating the dynamic exponents from the effective static exponents, $\rho_{eff}$ and $\nu_{eff}$, shown in Fig. 3. As shown in the right column of Fig. 4 (green crosses), there is also good agreement between the predicted and the observed ring weight dependence of the asymptotic CM diffusion constants.
4.3.3 Brownian dynamics of randomly quenched trees
@@
The emerging dynamics resulting from the exclusive application of tree node moves to our double-folded rings is presented in the central column of Fig. 6. (Detailed analysis is provided in Figs. S9–S11 in Sec. VIII of the Supplementary Material).
@@
For ideal and self-avoiding trees, our data are in good agreement with the predictions of the Rouse model (Tabel 1). The ring CM essentially exhibits simple diffusion, $g_{3}(t)\sim t/N_{m}$, with a diffusion constant inversely proportional to the number of ring monomers. A slowdown occurs independently of the ring size only over the first $50$ MC sweeps in the ideal trees, while in the self-avoiding case, the slowdown persists for a longer duration, up to $1000$ sweeps. The monomer dynamics, $g_{1}(t)$, in the intermediate intra-chain scaling regime also approaches the prediction $g_{1}(t)\sim t^{\frac{2\nu}{2\nu+1}}$ of the Rouse model (Table 1). Again, the slow decay of the effective exponents reported in the detailed analysis in Fig. 6 agrees, at least qualitatively, with the observed crossover of the static exponents in Fig. 3.
@@
The situation is more complicated in the melt case. While the observed monomer MSDs are again in reasonable accord with the prediction $g_{1}(t)\sim t^{0.4}$ of the Rouse model, the ring CM clearly exhibits anomalous diffusion in the intermediate regime. This translates to a scaling of $D\sim N^{-1.6\pm 0.1}$ over the present range of ring sizes and thus a faster decay than expected from the Rouse model (right column in Fig. 4). We note that experimental studies on ring polymers in the melt state have also reported a substantial slowdown in COM dynamics  58, 25, 31, which occurs before the system reaches the free diffusion regime.
4.3.4 Combining monomer and node moves
@@
Given that the node dynamics is significantly faster than the monomer dynamics in ideal and isolated self-avoiding systems (see Fig. 4), it is not surprising that the former dominates the combined dynamics entirely, causing the resulting scaling plots to essentially collapse onto those in the central column.
It is tempting to draw similar conclusions in the melt case, but with respect to the left column of Fig. 6 and a faster RiAO-like dynamics (bottom row in Fig. 4). Yet, the situation is more subtle since the overal dynamics is accelerated with respect to the pure dynamics of either monomer moves or tree node moves. The excellent scaling of the data reported in the bottom right panel of Fig. 6 is thus in itself an interesting observation. In both figures we have tentatively included slopes predicted by the FLG model, but it is difficult to draw any quantiative conclusions at this point.
V.  DISCUSSION
@@
The remaining challenge is to develop a protocol for the analysis of data obtained in mixed dynamical schemes. In particular, this protocol should allow us to identify and distinguish situations where the contributions of different local modes to the emergent dynamics are independent from those where the emergent dynamics arise from coupling effects.
@@
In section IV in the supplementary material, we propose, in Eq. (S6), an approximate plausible generalization of Eq. (11) for the full time dependence of the null model, for the monomer and CM MSDs in a mixed scheme. In Fig. 6 in section V in the supplementary material, we show that Eq. (S6) works well in a case where we expect the null model to apply: the dynamics emerging from the combination of repton and tree node moves, which are indeed completely decoupled within our model.
@@
While the ergodic RiAO dynamics provides an interesting example of the emergence of a qualitatively different dynamics from the combination of local moves (see section VI in the supplementary material), in the remainder of the (main) text, we focus on the emerging dynamics resulting from the combination of monomer and node moves in different proportions, $0\leq p_{m}\leq 1$, see Eq. 10. As evident from Figure 4 (in the main text) and Figure S7 (in the supplementary material), in interacting systems, this combination can lead to an accelerated dynamics compared to the two limiting cases: the RiAO dynamics ($p_{m}=1$) and the dynamics of quenched trees ($p_{m}=0$). This raises the question of how this acceleration depends on the ring weight, $N_{m}$, and the mixing parameter, $p_{m}$.
@@
For our mixing rule, Eq. (10), of the monomer and node moves, the prediction of the null model, Eq. (11), reads $D(p_{m})=p_{m}D_{RiAO}+(1-p_{m})D_{\mathrm{QTree}}$ or, equivalently, $D(p_{m})/D_{\mathrm{QTree}}=1+(D_{RiAO}/D_{\mathrm{QTree}}-1)p_{m}$. The panels in the left column of Fig. 7 show the corresponding results for the three types of investigated systems (ideal rings, self-avoided rings and rinds in melts). Note that the normalization to $D_{\mathrm{QTree}}$ allows us to show results for rings of different sizes on a common linear scale (Raw results of $D$ are presented in Fig. S17 and in Table. II of the Supplementary Material.).
@@
Results reveal the importance and character of deviations of the emergent dynamics from the null model. Namely, while the emergent dynamics in the melt is significantly accelerated and dominated by the deviations from the null model, the corrections appear to be more of a qualitative nature for self-avoiding rings and statistically insignificant for ideal rings. This said, our decision to normalize the diffusion constants in all three cases by those of the asymptotically faster Brownian Tree Dynamics (right column of Fig. 4) can lead to the wrong impression that $D(p_{m})$ extrapolates to zero for finite ring sizes (as it does indeed in the case of the slithering dynamics in Fig. S6(d) in the supplementary material). However, in the present case $D(p_{m})/D_{\mathrm{QTree}}$ remains finite for all $N_{m}$, albeit much smaller than unity for ideal and self-avoiding double-folded rings and equal to zero only in the limit of infinite chain length.
In the right column of Fig. 7, we show the difference
$\Delta D(p_{m})=D(p_{m})-\left(p_{m}D_{\mathrm{RiAO}}+(1-p_{m})D_{\mathrm{QTree}}\right)$ (12)
@@
between the observed asymptotic CM diffusion constant and the expected value from the null model. To simplify a comparison with the data shown in the panels in the left column, we do not show absolute values for $\Delta D(p_{m})$ but $\Delta D(p_{m})/D_{\mathrm{QTree}}$. If $\Delta D(p_{m})$ vanishes by construction in the two limits $p_{m}=0$ and $p_{m}=1$ (i.e., a value of zero is in this case a “true” and not just an apparent zero), our data suggest an intriguing symmetry in how the node and the monomer dynamics combine.
Figure 7: Left column: The asymptotic center-of-mass (CM) diffusion constants for different systems, obtained using a combination of monomer and node moves in varying proportions ($0\leq p_{m}\leq 1$), as described in Eq. 10. The dashed lines represent the predictions from the null model (Eq. 11). Right column: The difference between the observed asymptotic CM diffusion constants and the expected values from the null model, as $\Delta D(p_{m})=D(p_{m})-\left(p_{m}D_{\mathrm{RiAO}}+(1-p_{m})D_{\mathrm{QTree}}\right)$. For better visualization across different system sizes, all data are normalized by $D_{\text{QTree}}(N_{m})$.
@@
Interestingly, while contributions to the ring CM diffusion constant from two coupled A-type and B-type dynamics may take multiple forms (see Sec. VII in the supplementary material), our results for $\Delta D_{AB}$ reported in the right column of Fig. 7 suggest that the coupling contribution adopts the simplest possible form of
@@
Is this (in hindsight) obvious and inevitable or, alternatively, does our observation constrain the choice of admissible microscopic coupling mechanisms? Since the CM diffusion of our rings is diffusive in the long time limit, we consider as the simplest model the coupling of two diffusion steps, $\vec{\delta}_{A}$ and $\vec{\delta}_{B}$, for a particle moving under the influence of two random processes $A$ and $B$ such that $\left\langle\left|\vec{\delta}_{A}\right|^{2}\right\rangle\propto p_{A}D_{A}$ and $\left\langle\left|\vec{\delta}_{B}\right|^{2}\right\rangle\propto p_{B}D_{B}$. To define the additional displacement, $\vec{\delta}_{AB}=\delta_{AB}\hat{\delta}_{AB}$, due to the coupling of the two steps, we need to specify its magnitude $\delta_{AB}\equiv\lVert\vec{\delta}_{AB}\rVert$ and direction, $\hat{\delta}_{AB}\equiv\vec{\delta}_{AB}/\lVert\vec{\delta}_{AB}\rVert$. For the magnitude, admissible combinations of $\delta_{A}$ and $\delta_{B}$ have to obey the same dimensionality constraint, that we have discussed above for the diffusion constant. With simple vector operations like $\lVert\vec{\delta}_{A}\times\vec{\delta}_{B}\rVert$ and $\vec{\delta}_{A}\cdot\vec{\delta}_{B}$ having the unit of the square of a length, one arrives rather naturally at a scaling, $\delta_{AB}\propto\sqrt{\delta_{A}\delta_{B}}$, of the length of the coupling steps with the geometric mean of the elementary step lengths, that may be modulated by geometric factors. On first sight, this seems to directly lead to the observed scaling $\Delta D_{AB}=\langle\delta_{AB}^{2}\rangle\propto\sqrt{\langle\delta_{A}^{2}\rangle\,\langle\delta_{B}^{2}\rangle}\propto\sqrt{p_{A}D_{A}p_{B}D_{B}}$ of the coupling contribution to the diffusion constant. However, the relation $\Delta D_{AB}\equiv D_{AB}-p_{A}D_{A}-p_{B}D_{B}$ only holds, if additionally $\hat{\delta}_{AB}=\left(\vec{\delta}_{A}\times\vec{\delta}_{B}\right)/\lVert\vec{\delta}_{A}\times\vec{\delta}_{B}\rVert$ or if the direction of the coupling steps, $\hat{\delta}_{AB}$, is not correlated with $\vec{\delta}_{A}$ and $\vec{\delta}_{B}$. In contrast, if $\hat{\delta}_{AB}$ lies in the plane spanned by the elementary steps, $\vec{\delta}_{A}$ and $\vec{\delta}_{B}$, correlations lead to more complicated relations of the general type Eq. (S13) in the supplementary material.
As our final point, we address the extrapolation of our data to ring sizes beyond the range studied in the present work. On first sight, the diffusion constants reported in Fig.4 seem compatible with three power laws for the two pure and the mixed dynamics with $p_{m}=0.5$. In particular, the data for the total diffusion constant in the mixed case are close to the prediction of the FLG model 37. However, this interpretation is incompatible with our observations for the full range of $p_{m}$ values shown in Fig.7. In particular, for other values of $p_{m}$ our data for $D$ do not follow a uniform power law (see Fig. S16 in the Supplementary Material). Instead, our analysis suggests that the total diffusion constant for the mixed dynamics can be expressed as:
@@
where
$\delta D_{AB}(N)\equiv\frac{\Delta D_{AB}(N)}{\sqrt{p_{A}p_{B}}}=a(N)\sqrt{D_{A}(N)D_{B}(N)}$ (15)
is the magnitude of the contribution to the diffusion constant arising from the coupling.
In spite of the vastly different values of $D_{\mathrm{QTree}}$ for isolated self-avoiding rings and for rings in a melt being approximately two orders of magnitude smaller (As shown in Fig.4 (main text) and in Table II (Supplementary Material)), Fig. 8 shows that the absolute magnitude of the $\delta D$ arising from the coupling with the RiAO-like dynamics is quite similar for the two systems. Moreover, in both cases $\delta D$ appears to be of the order of $D_{\mathrm{QTree}}$, i.e. the dimensionless prefactor $a(N)$ in Eq. (15) for the ring-size dependent efficiency of the coupling is by no means a trivial constant. For isolated self-avoiding rings $\lim_{N\rightarrow\infty}\delta D/D_{\mathrm{QTree}}>0.4$ and our data are in good agreement with a Rouse-like scaling, $\delta D\sim N^{-1}$. Our data for melts are less clear. While they show no sign of converging to the FLG prediction of $\delta D\sim N^{-5/3}$, they are compatible with a slow crossover to a Rouse-like scaling. In particular, $\lim_{N\rightarrow\infty}\delta D/D_{\mathrm{QTree}}>12$. Notably the total CM diffusion constant in the melt case seems to be dominated asymptotically by the mobility arising from the coupling of the Brownian tree and the RiAO-like dynamics.
VI.  SUMMARY AND CONCLUSION